OpenStudy (anonymous):

(4x^-2 y^5)^-2 (2xy^4)^3 I no the answer is 1/2 x^7 y^2 but I need to show my work?

OpenStudy (whpalmer4):

You know the answer, but you can't show your work? Did you do the problem, or find the answer elsewhere? \[(4x^{-2}y^5)^{-2}(2xy^4)^3\] Some basic properties of exponents that you can use to simplify this (and show your work :-) \[(ab)^n = a^nb^n\]\[a^na^m = a^{n+m}\]\[(a^n)^m = a^{n*m}\]\[a^{-n} = \frac{1}{a^n}\]